When serving a tennis ball, a player hits the ball when its velocity is zero (at the highest point of a vertical toss). The racquet exerts a force of 540 N on the ball for 5.00 milliseconds, giving it a final velocity of 45.0 m/s. Find the mass of the ball.



1
Expert's answer
2021-03-08T08:29:24-0500
"F\Delta t=m\Delta v,""F\Delta t=mv_f-mv_i=mv_f,""m=\dfrac{F\Delta t}{v_f}=\dfrac{540\ N\cdot5\cdot10^{-3}\ s}{45\ \dfrac{m}{s}}=0.06\ kg=60\ g."